# Modeling the Hive Mind: Aligning Large Language Models with Online Communities

### v1

@@ Abstract
Large language models (LLMs) have shown promise in representing individuals and communities, offering new ways to study complex social dynamics. However, effectively aligning LLMs with specific human groups and systematically assessing the fidelity of the alignment remains a challenge. This paper presents a robust framework for aligning LLMs with online communities via instruction-tuning and comprehensively evaluating alignment across various aspects of language, including authenticity, emotional tone, toxicity, and harm. We demonstrate the utility of our approach by applying it to online communities centered on dieting and body image. We administer an eating disorder psychometric test to the aligned LLMs to reveal unhealthy beliefs and successfully differentiate communities with varying levels of eating disorder risk. Our results highlight the potential of LLMs in automated moderation and broader applications in public health and social science research [^1] Our data and code will be available upon publication..
@@ 1 Introduction
Large language models (LLMs) have demonstrated an unparalleled ability to generate detailed, nuanced responses to natural language prompts, suggesting potential for their use in creating high-fidelity proxies of people Simmons and Hare (2023). Leveraging LLMs to create digital representations of individuals and human groups could provide powerful tools for studying psycho-social dynamics of human behavior, enhancing and personalizing human-computer interactions, and moderating online spaces to promote prosociality, enhance safety, and reduce harm.
@@
To create digital representations of human subgroups, researchers have aligned LLMs to subgroups via steering—i.e., instructing the LLM to mimic the target subgroup by specifying its core characteristics in the prompt Santurkar et al. (2023); Durmus et al. (2023). However, this approach does not solve LLMs’ misalignment with the target subgroup. Other methods for aligning LLMs to human subgroups include finetuning the base LLMs [^2] By “base LLMs” we refer to LLMs that are not finetuned to follow instructions like GPT-2 on data generated by specific subgroups Jiang et al. (2022b); He et al. (2024c). Although this method can produce models that reflect the linguistic patterns of the target population, these finetuned models often lack the flexibility to follow diverse instructions, limiting their utility.
@@
Another key challenge in developing digital representations of human subgroups is evaluating the alignment between the LLM and the target group. Traditional methods compare the LLM’s responses to surveys with those of the target group Santurkar et al. (2023); Durmus et al. (2023), but this approach misses critical aspects of human expression like emotional reactions He et al. (2024b). Additionally, surveys are not scalable due to their cost and time requirements, particularly for marginalized or hard-to-reach groups. Besides, mapping organically-formed online communities to clear demographic identities greatly complicates alignment evaluation.
@@
To address these challenges, we propose a framework for aligning LLMs with online communities through instruction-tuning in a fully unsupervised manner. Additionally, we introduce a comprehensive evaluation framework to assess alignment. This enables the creation of high-fidelity digital representations of online communities, paving the way for new research into human behavior, content moderation, public mental health, and social science. As one example, we can administer psychometric instruments to these digital proxies to identify at-risk communities prone to psychopathologies.
@@
Our alignment method takes a corpus of social media posts (e.g., tweets) from an online community and creates a set of demonstrations (instruction-response pairs). In each demonstration (Figure 1), the instruction specifies the task (e.g., generate a tweet) with the response being the exact tweet. We then finetune an LLM on these demonstrations to align it with the community. To assess alignment, we generate a synthetic text corpus using the finetuned LLM and compare it to the original posts along four key aspects: 1) authenticity, 2) emotional tone, 3) toxicity, and 4) harm. These dimensions capture the essential features of online social communication, ensuring the aligned LLM accurately reflects the semantics, affect, and style of the target group’s discourse.
@@
To demonstrate our framework’s utility, we analyze Twitter discussions in diet and fitness communities, where harmful attitudes about body image exist. While these communities can offer support and encouragement, they often promote unhealthy behaviors and beliefs that put people at risk for developing eating disorders (EDs). Applying traditional psychometric instruments to screen individuals for EDs is impractical and potentially unethical; instead, we use our framework to align LLMs with these communities through automatically generated demonstrations and evaluate alignment to show that the finetuned LLMs outperform baseline LLMs in creating high-fidelity proxies of online communities. We then apply an ED screening questionnaire to community-aligned LLMs, revealing significant differences between communities: pro-anorexia communities show a high risk of unhealthy behaviors, while those critical of the diet culture exhibit the lowest risk. These findings highlight our framework’s potential for automated moderation by distinguishing communities with varying levels of ED risk.
Our framework offers a scalable approach to modeling and analyzing online communities, with broad implications for understanding and mitigating harmful behaviors. By applying this method to ED communities, we demonstrate its potential to contribute to public health and social science research, highlighting the value of LLMs in studying complex social dynamics.
@@ 2 Related Work
Aligning LLMs to Subgroups
There is growing literature Simmons and Hare (2023) on aligning LLMs to diverse human subgroups to mimic their language and mindsets. Researchers have aligned LLMs by steering them towards particular demographic groups Santurkar et al. (2023); Durmus et al. (2023); He et al. (2024b), e.g., by including the target subgroup in the prompt. However, their findings reveal that steering does not solve the model’s misalignment with the target subgroup. Moreover, it is non-trivial to summarize an organically-formed community (e.g., communities in retweet networks) into a concise description that can be used in steering.
Others have aligned LLMs with different subgroups by finetuning the model on the text generated by the subgroups. Jiang et al. (2022b) propose communityLM by finetuning two GPT-2 models Radford et al. (2020) using causal language modeling on tweets from liberals and conservatives, and probing their worldviews from their corresponding finetuned models. He et al. (2024c) extend communityLM to probe the views of organically-formed online communities and make use of the interactions between different communities. However, GPT-2 is not instruction-tuned and is not able to answer questions in various formats, like the psychometric instruments we discuss in §$\S$6. He et al. (2024a) use an advanced LLM (e.g., Claude-3) to distill knowledge from the community’s raw data and generate high-quality instruction-response pairs, where the instructions aim to query the community’ mindset, and the corresponding responses are abstracted from the ideas conveyed in the raw data. The generated instruction-response pairs are used to finetune a foundational LLM (e.g., Llama-3) for alignment. However, the API costs of querying the advanced LLM are non-negligible.
@@
Existing works Santurkar et al. (2023); Durmus et al. (2023) measure an LLM’s alignment with a target subgroup using multi-choice surveys. Specifically, they prompt the LLM to respond to a survey question from the perspective of a subgroup and then compare the LLM-generated distribution over the different options of the question to that of the survey respondents belonging to the target group. However, collecting survey responses can be costly and time-consuming. Also, responses on sensitive topics, such as mental health, may be biased due to stigma and social desirability bias Gordon (1987). Our framework evaluates LLM alignment by comparing the LLM-generated synthetic text to the original text written by humans is significantly more scalable, unbiased, and cost-effective.
@@
LLMs can respond to psychometric instruments that were originally designed to assess individual human psychological and emotional states. Researchers have administered these instruments to LLMs to probe their decision-making processes, reasoning abilities, cognitive biases, and other psychological traits—Pellert et al. (2024) call this practice “AI Psychometrics”. Coda-Forno et al. (2023) show that GPT-3.5 generated consistently high scores on responses to a widely used anxiety questionnaire. Tanmay et al. (2023) measure GPT-4’s moral reasoning abilities by applying an ethical measurement instrument for individuals. Researchers also administer personality tests to LLMs to identify their personality traits Jiang et al. (2022a); Lu et al. (2023); Serapio-García et al. (2023). In contrast, we apply psychometric questionnaires to a specific online community—via a finetuned LLM—to learn more about the mindset of the community members. We show that this helps reveal unhealthy beliefs within these communities and even identify pathologies, like harmful cognitions associated with EDs.
@@
Previous research has focused on identifying harmful content and at-risk users within these communities. For example, Chancellor et al. (2016a) developed a lexical classifier to predict posts moderated by Instagram for self-harm content, comparing pro-recovery and pro-ED communities Chancellor et al. (2016b, c). In contrast, our study examines the collective mindset of these communities as expressed through their discussions, using advanced language models to assess attitudes toward mental health and body image issues.
@@ 3.1 Data Collection
We collected 2.6M tweets from 557K users from October 2022 to March 2023 using ED-related keywords to query Twitter. For keywords, we start with a set of terms that promote ED  Chancellor et al. (2016a); Pater et al. (2016), such as thinspo (thin inspiration), proana (pro-anorexia), and promia (pro-bulimia), among others. We remove spam terms yielding unrelated content, such as skinny. We expanded the query set to include closely related topics such as diet and weight loss through terms such as (ketodiet, weightloss, $\ldots$), and anti-diet culture (bodypositivity, dietculture, $\ldots$). See Appendix A.1 for the full set of search terms.
@@ 3.2 Identifying Communities
Figure 2: The framework of our method. (1) We align an LLM (Llama-3) to an online community by finetuning the LLM to follow instructions on the task of generating tweets written by users in the community. (2) To prove the effectiveness of alignment, we compare three tweet corpora for each community: human-written tweets $D_{i}$, RAG LLM-generated tweets $D^{RAG}_{i}$, and finetuned LLM-generated tweets $D^{ft}_{i}$. We show that $D^{ft}_{i}$ is closer to $D_{i}$ than $D^{RAG}_{i}$ is, along the following aspects: (a) A classifier trained to classify the tweet origin (what community the tweet belongs to) on $\mathbb{D}=\{D_{i}\}_{i=1}^{n}$ performs equally well on $\mathbb{D}^{ft}=\{D^{ft}_{i}\}_{i=1}^{n}$, but not on $\mathbb{D}^{RAG}=\{D^{RAG}_{i}\}_{i=1}^{n}$; (b) the emotion and toxicity distributions of $D^{ft}_{i}$ are much closer to that of $D_{i}$ compared to $D^{RAG}_{i}$; (c) the semantic embeddings of $D^{ft}_{i}$ are closer to that of $D_{i}$ in the embedding space than that of $D^{RAG}_{i}$ are; (d) a human annotator decides that $D^{ft}_{i}$ is more aligned to the underlying distribution of $D_{i}$ than $D^{RAG}_{i}$ is; (e) two ED experts determine that $D^{ft}_{i}$ carries harmful narratives that are more similar to $D_{i}$ than $D^{RAG}_{i}$. (3) As the LLM is aligned with the community and can speak in the voice of that community, we administer an ED questionnaire to screen the community for EDs.
@@ 4.1 Constructing Instruction-Response Pairs
For each community $C_{i}$, we use tweets in $D_{i}$ as the responses verbatim in the demonstrations. To create instructions that can be answered by the tweets, we focus on the tweet generation task. We curate an instruction pool of 20 different instruction templates (Table 5 in Appendix B.1). For a community, a tweet is paired with an instruction randomly sampled from the instruction pool. As a result, the community has a maximum of 10K demonstrations $Z_{i}=\{(x_{j},y_{j})\}_{j=1}^{m}$ for tweet generation, where $m$ is the size of the community’s text corpus $D$.
@@ 4.2 Instruction Tuning LLMs
For each community $C_{i}$, we align a Llama-3 model $f^{\prime}_{i}$ AI@Meta (2024) to the community using its demonstration corpus $Z_{i}$. The LLM is finetuned on 4 Tesla H100-80GB GPUs with batch size 8 for 3 epochs, which takes about 3 hours.
@@ 5.1 Synthetic Corpus Generation
Given a community $C_{i}$, we create a synthetic corpus $D^{ft}_{i}$ by prompting an LLM $f^{\prime}_{i}$ aligned to the community to generate tweets. To diversify the LLM generations, we compile a set of 27 topics relevant to ED discussions, such as thinspo, fitspo, and bonespo (Appendix C.1), and prompt LLMs to generate tweets on these topics. When generating tweets on a topic, we reuse the diverse instructions from the instruction pool (Table 5 in Appendix B.1). An example instruction is “What would you tweet about fasting?” For each topic, the LLM generates 400 tweets, resulting in a synthetic corpus $D^{ft}_{i}$ with 10,800 tweets for all 27 topics.
We detail the number of tweets in the community’s original text corpus $D_{i}$ that contain the keyword(s) of each of the 27 topics in Table 7 in the Appendix C.1. We observe that $D_{i}$ contains a very limited number of tweets discussing these topics. This is because we removed the hashtags in tweet processing, and these keywords usually appear in the hashtags. Consequently, when the LLM is finetuned on $D_{i}$, it is not extensively exposed to tweets that are directly related to these topics. This ensures that the synthetic corpus $D^{ft}_{i}$ does not simply replicate $D_{i}$. Instead, when the finetuned LLM $f^{\prime}_{i}$ generates synthetic tweets on these 27 topics, it extrapolates from existing data in $D_{i}$ to predict how community members might discuss these previously unseen topics.
Baseline We use the LLM with retrieval-augmented generation (RAG) Lewis et al. (2020) as a baseline. We do not finetune the RAG model. For a community $C_{i}$, when prompting the model to generate synthetic tweets on topic $t$, we retrieve 250 tweets, consisting of (1) the tweets containing the topic keyword(s), if available, and (2) randomly sampled tweets from $D_{i}$. Each retrieved tweet is truncated at 20 tokens. We include the retrieved tweets in the prompt, instruct the model to learn the community’s mindset from the tweets, and generate synthetic tweets. See Appendix C.2 for the complete prompting template. The synthetic corpus on all topics from the RAG model is denoted as $D^{RAG}_{i}$.

### v2

@@ Abstract
Large language models (LLMs) have shown promise in representing individuals and communities, offering new ways to study complex social dynamics. However, effectively aligning LLMs with specific human groups and systematically assessing the fidelity of the alignment remains a challenge. This paper presents a robust framework for aligning LLMs with online communities via instruction-tuning and comprehensively evaluating alignment across various aspects of language, including authenticity, emotional tone, toxicity, and harm. We demonstrate the utility of our approach by applying it to online communities centered on dieting and body image. We administer an eating disorder psychometric test to the aligned LLMs to reveal unhealthy beliefs and successfully differentiate communities with varying levels of eating disorder risk. Our results highlight the potential of LLMs in automated moderation and broader applications in public health and social science research  [^1] Our data and code are available at https://github.com/Davidchu11381/llm_align_eval. .
@@ 1 Introduction
Large language models (LLMs) have demonstrated an exceptional ability to generate nuanced responses to natural language prompts, suggesting their potential for creating high-fidelity proxies of people Simmons and Hare (2023). Digital representations of human groups (digital twins) are computational models that mimic collective behaviors, social interactions, and communication patterns in real-world communities Rossetti et al. (2024b). Leveraging LLMs to build these representations offers powerful tools for studying human behavior, enhancing human-computer interactions, and moderating online spaces to foster prosociality and safety.
@@
To create such digital twins, researchers align LLMs to subgroups through steering—prompting the LLM to mimic the target subgroup with its key characteristics Santurkar et al. (2023); Durmus et al. (2023). However, this approach does not fully resolve the misalignment between LLMs and the target subgroup. Another method is finetuning base LLMs [^2] By “base LLMs” we refer to models not finetuned for instruction following, such as GPT-2, on data from specific subgroups Jiang et al. (2022b); He et al. (2024d). While this can produce models reflecting subgroup linguistic patterns, these finetuned models often lack the flexibility to follow diverse instructions, limiting their broader applicability.
@@
Another key challenge in developing digital representations of human subgroups is evaluating the alignment between the LLM and the target group. Traditional methods compare the LLM’s responses to surveys with those of the target group Santurkar et al. (2023); Durmus et al. (2023), but this approach misses critical aspects of human expression like emotional reactions He et al. (2024b). Additionally, surveys are not scalable due to their cost and time requirements, particularly for marginalized or hard-to-reach groups. In addition, mapping diverse online communities to clear demographic identities greatly complicates alignment evaluation.
@@
To address these challenges, we propose a framework for aligning LLMs with online communities on instructions that are created in a fully unsupervised manner. Additionally, we introduce a comprehensive evaluation framework to assess alignment. This enables the creation of high-fidelity digital representations of online communities, paving the way for new research into human behavior Jiang et al. (2022c), content moderation He et al. (2024c), public mental health Sharma et al. (2024), and social science Grossmann et al. (2023). As one example, we can administer psychometric instruments to these digital replicas to identify at-risk communities prone to psychopathologies.
@@
Specifically, our alignment method takes a corpus of social media posts (e.g., tweets) from an online community and creates a set of demonstrations (instruction-response pairs) based on the posts. In each demonstration, as shown in Figure 1, the instruction specifies the task (e.g., tweet generation) with the response being the exact tweet. We then finetune an LLM on these demonstrations to align it with the community. To assess alignment, we generate a synthetic text corpus using the finetuned LLM and compare it to the original posts along four key aspects: 1) authenticity, 2) emotional tone, 3) toxicity, and 4) harm. These dimensions capture the essential features of online social communication, ensuring the aligned LLM accurately reflects the semantics, affect, and style of the target group’s discourse.
@@
To demonstrate our framework’s utility, we analyze Twitter discussions in diet and body image communities, where harmful body image attitudes persist. Twitter’s emphasis on user connection and lax content moderation allows communities to organically form and freely express their voice. While these communities can offer support and encouragement, they often promote unhealthy behaviors and normalize beliefs that put people at risk for developing eating disorders (EDs). Applying traditional psychometric instruments to screen individuals in those online spaces for EDs is impractical and potentially unethical (see Ethics Statement). Instead, we use our framework to align LLMs with these communities through automatically generated demonstrations and evaluate alignment to show that the finetuned LLMs outperform baseline LLMs in creating high-fidelity proxies of online communities. We then apply an ED screening questionnaire to community-aligned LLMs, revealing significant differences between communities: pro-anorexia communities show a high risk of unhealthy behaviors, while those critical of the diet culture exhibit the lowest risk. These findings highlight our framework’s potential for automated moderation by distinguishing communities with varying levels of ED risk.
Our framework offers a scalable approach to modeling and analyzing online communities, with broad implications for understanding and mitigating harmful behaviors. By applying this method to diet and body image communities, we demonstrate its potential to contribute to public health and social science research, highlighting the value of LLMs in studying complex social dynamics.
@@ 2 Related Work
Digital Representations of Human Subgroups
Digital twins—precise virtual replicas of complex real-world systems—are increasingly employed for advanced analysis and experimentation Tao and Qi (2019); Grieves (2011), particularly in monitoring human behaviors and health outcomes Ferdousi et al. (2022); Shengli (2021); El Saddik et al. (2019). Recent advances leverage social media data to provide deeper insights into human interactions Olad and Valilai (2020), exemplified by Rossetti et al. (2024a)’s Y social, which uses LLMs to simulate social media interactions and study network dynamics in controlled environments.
Building on this foundation, researchers have explored various approaches to align LLMs with diverse human subgroups Simmons and Hare (2023). While some have attempted prompt-based steering towards specific demographic groups Santurkar et al. (2023); Durmus et al. (2023); He et al. (2024b), this approach has shown limitations in achieving true alignment, particularly with organically-formed communities. More promising results have come from finetuning approaches: Jiang et al. (2022b) developed communityLM by finetuning GPT-2 models Radford et al. (2019) on politically divergent tweets, while He et al. (2024d) extended this to examine broader community interactions. Most recently, He et al. (2024a) proposed using advanced LLMs to distill community knowledge into instruction-response pairs for finetuning, though this approach faces cost constraints.
Our study aims to advance this field by developing a framework that uses LLMs to create digital representations of online communities, specifically focusing on analyzing collective mental well-being.
@@
Existing works Santurkar et al. (2023); Durmus et al. (2023) measure an LLM’s alignment with a target subgroup using multi-choice surveys. Specifically, they prompt the LLM to respond to a survey question from the perspective of a subgroup and then compare the LLM-generated distribution over the different options of the question to that of the survey respondents belonging to the target group. However, collecting survey responses can be costly and time-consuming. Also, responses on sensitive topics, such as mental health, may be biased due to stigma and social desirability bias Gordon (1987). Our framework that evaluates LLM alignment by comparing the LLM-generated synthetic text to the original text written by humans is significantly more scalable, unbiased, and cost-effective.
@@
LLMs can respond to psychometric instruments designed for humans, with researchers using these tools to examine LLMs’ decision-making, reasoning, and cognitive traits—a practice termed “AI Psychometrics” Pellert et al. (2024). Research shows LLMs can engage with various psychometric tools, from anxiety questionnaires Coda-Forno et al. (2023) to moral reasoning assessments Tanmay et al. (2023) and personality tests Jiang et al. (2022a); Lu et al. (2023); Serapio-García et al. (2023). Our work differs by using these instruments via a finetuned LLM to analyze specific online communities, helping identify unhealthy beliefs and potential pathologies like eating disorders-related cognitions.
@@
Previous research has focused on identifying harmful content and at-risk users within these communities. For example, Chancellor et al. (2016a) develop a lexical classifier to predict posts moderated by Instagram for self-harm content, comparing pro-recovery and pro-ED communities Chancellor et al. (2016b, c). In contrast, our study examines the collective mindset of these communities as expressed through their discussions, using advanced language models to assess attitudes toward mental health and body image issues.
@@ 3.1 Data Collection
We collected 2.6M tweets from 557K users from October 2022 to March 2023 using ED-related keywords to query Twitter. For keywords, we start with a set of terms that promote ED  Chancellor et al. (2016a); Pater et al. (2016), such as thinspo (thin inspiration), proana (pro-anorexia), and promia (pro-bulimia), among others. We remove spam terms yielding unrelated content, such as skinny. We expanded the query set to include closely related topics such as diet and weight loss through terms such as (ketodiet, weightloss, $\ldots$), and anti-diet culture (bodypositivity, dietculture, $\ldots$). See Appendix A.1 for the full set of search terms and why we selected them.
@@ 3.2 Identifying Communities
Figure 2: The framework of our method. (1) We align an LLM (Llama-3) to an online community by finetuning the LLM to follow instructions on the task of generating tweets written by users in the community. (2) To prove the effectiveness of alignment, we compare three tweet corpora for each community: human-written tweets $D_{i}$, LLM-Context-generated tweets $D^{context}_{i}$, and finetuned LLM-generated tweets $D^{ft}_{i}$. We show that $D^{ft}_{i}$ is closer to $D_{i}$ than $D^{context}_{i}$ is, along the following aspects: (a) A classifier trained to classify the tweet origin (what community the tweet belongs to) on $\mathbb{D}=\{D_{i}\}_{i=1}^{n}$ performs better on $\mathbb{D}^{ft}=\{D^{ft}_{i}\}_{i=1}^{n}$, than on $\mathbb{D}^{context}=\{D^{context}_{i}\}_{i=1}^{n}$; (b) the emotion and toxicity distributions of $D^{ft}_{i}$ are much closer to that of $D_{i}$ compared to $D^{context}_{i}$; (c) the semantic embeddings of $D^{ft}_{i}$ are closer to that of $D_{i}$ in the embedding space than that of $D^{context}_{i}$ are; (d) a human annotator decides that $D^{ft}_{i}$ is more aligned to the underlying distribution of $D_{i}$ than $D^{context}_{i}$ is; (e) two ED experts determine that $D^{ft}_{i}$ carries harmful narratives that are more similar to $D_{i}$ than $D^{context}_{i}$ does. (3) As the LLM is aligned with the community and can speak in the voice of that community, we administer an ED questionnaire to screen the community for EDs.
@@ 4.1 Constructing Instruction-Response Pairs
For each community $C_{i}$, we use tweets in $D_{i}$ as the responses verbatim in the demonstrations. To create instructions that can be answered by the tweets, we focus on the tweet generation task. We curate an instruction pool of 20 different instruction templates (Table 5 in Appendix B.1). We diversify the prompts to improve the model’s text-generation capabilities and enhance its robustness to linguistic variations Salinas and Morstatter (2024). For a community, a tweet is paired with an instruction randomly sampled from the instruction pool. As a result, the community has a maximum of 10K demonstrations $Z_{i}=\{(x_{j},y_{j})\}_{j=1}^{m}$ for tweet generation, where $m$ is the size of the community’s text corpus $D$.
@@ 4.2 Instruction Tuning LLMs
For each community $C_{i}$, we align a Llama-3 model $f^{\prime}_{i}$ Dubey et al. (2024) to the community using its demonstration corpus $Z_{i}$. The LLM is finetuned on 4 Tesla H100-80GB GPUs with batch size 8 for 3 epochs, which takes about 3 hours.
@@ 5.1 Synthetic Corpus Generation
Given a community $C_{i}$, we create a synthetic corpus $D^{ft}_{i}$ by prompting an LLM $f^{\prime}_{i}$ aligned to the community to generate tweets. To diversify the LLM generations on different topics, we compile a set of 27 topics relevant to ED discussions, such as thinspo, fitspo, and bonespo (Appendix C.1), and prompt LLMs to generate tweets on these topics. When generating tweets on a topic, we reuse the instructions from the instruction pool (Table 5 in Appendix B.1), with topic-oriented generation. An example instruction is “What would you tweet about fasting?” For each topic, the LLM initially generates 1000 tweets, resulting in a synthetic corpus $D^{ft}_{i}$ with 27,000 tweets for all 27 topics (see Appendix C.4 for examples). To encourage diversity in $D^{ft}_{i}$, we remove duplicate tweets. In addition, to ensure coherence, a generated tweet is filtered out if its perplexity score is above 400, as evaluated by BERTweet.
A natural concern is that $D^{ft}_{i}$ is simply a duplicate of $D_{i}$, as $f^{\prime}_{i}$ is finetuned on $D_{i}$. To this end, we detail the number of tweets in the community’s original text corpus $D_{i}$ that contain the keyword(s) for each of the 27 topics listed in Table 7 (see Appendix C.1). We observe that $D_{i}$ includes very few tweets discussing these topics because we eliminate hashtags during tweet processing, and these keywords typically appear in the hashtags. Consequently, when the LLM is finetuned on $D_{i}$, it is not extensively exposed to tweets directly related to these topics.
To further ensure that the synthetic corpus $D^{ft}_{i}$ does not simply replicate $D_{i}$, we omit generated tweets that are substantially syntactically similar to the human-written ones. Specifically, a tweet is removed from $D^{ft}_{i}$, if its ROUGLE-L similarity to any existing tweet in $D_{i}$ is greater than 0.7. As a result, when inspecting the synthetic corpus $D^{ft}_{i}$, we are essentially examining if the finetuned model $f^{\prime}_{i}$ is able to extrapolate from existing data in $D_{i}$ and predict how community members might discuss these previously unseen topics.
Finally, to ensure class balance, we sample 6000 generated tweets from each community. More details are provided in Appendix C.4.
Baseline We use the LLM with in-context learning (LLM-Context) as a baseline. We do not finetune this baseline model. For a community $C_{i}$, when prompting the model to generate synthetic tweets on topic $t$, we retrieve 250 tweets from $D_{i}$ as in-context examples, consisting of (1) the tweets containing the topic keyword(s), if available, and (2) randomly sampled tweets from $D_{i}$. Each retrieved tweet is truncated at 20 tokens. We include the retrieved tweets in the prompt, instruct the model to learn the community’s mindset from the tweets, and generate synthetic tweets. See Appendix C.2 for the complete prompting template. The synthetic corpus from LLM-Context is denoted as $D^{context}_{i}$.